OpenStudy (anonymous):

Evaluate the series 4-2+1-0.5+0.25 to S10. round to the nearest hundredth? Help me .....

OpenStudy (anonymous):

a=4. \[r=-\frac{ 2 }{ 4 }=-\frac{ 1 }{ 2 }\] \[S _{n}=\frac{ a \left( 1-r^n \right) }{ 1-r },put~ n=10\]